# CLQCD Collaboration

### v2

@@ IV Summary
In this work, we determine the charm quark mass using the $D_{s}$ meson mass and convert it to the $\overline{\textrm{MS}}$ scheme, and then predict the charmed meson masses and their decay constants. The results are summarized in Table 3. The values from PDG Workman et al. (2022), FLAG Aoki et al. (2022) and HPQCD Hatton et al. (2020a); Donald et al. (2014) are also listed there for comparison. All our predictions of the charmed meson masses are consistent with the experimental value at 0.3% level uncertainty. We also verified the unitarity of the CKM matrix elements involving the charm quark at 3% level, updated the constraints on the leptonic decay width of $D^{*}_{(s)}$, and provide the decay constants of the P-wave charmonium decay constants in the continuum with physical quark masses.
In the above predictions, we used the interpolated strange quark mass, which corresponds to the ”physical” $\eta_{s}$ mass Borsanyi et al. (2021). This mass may differ by 10-15% from the unitary strange quark mass on certain ensembles, and its use can effectively suppress the impact of the unphysical strange quark mass on those ensembles by an order of magnitude in the $D_{s}^{*}$ mass.
@@
Comparing to the overlap fermion where $Z_{V}=Z_{A}$ only depends on $m_{\rm PS}^{2}a^{2}$ with a tiny coefficient $\sim 0.02$ Wang et al. (2021); He et al. (2022), the clover fermion we used here has a much stronger $m_{\rm PS}^{2}a^{2}$ dependence with a coefficient $\sim 0.5$. It makes the continuum extrapolation using the heavy quark improved $Z^{c}_{V}$ normalization has much smaller discretization error for the charmed meosn matrix elements comparing to that using the mass-independent one. The values of the renormalization constants are collect in Table 2.
@@
In summary, we would expect the previous Zhang et al. (2022c); Xing et al. (2022); Liu et al. (2023, 2024a, 2024b); Meng et al. (2024); Han et al. (2024); Yan et al. (2024) and ongoing studies using the CLQCD ensembles Hu et al. (2024) for the charm physics can reach higher accuracy with the interpolated charm and strange quark masses and also the heavy quark improved normalization scheme presented in this work. More specifically, hopefully the systematic uncertainties coming from the unphysical light and strange quark masses and finite lattice spacing can be controlled using the CLQCD ensembles, in the spectrum, form factors, semi-letopic decay, and behaviors at the finite temperature and magnetic field of light and charmed meson and baryon.
@@ Acknowledgement
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu et al. (2024), which are generated on HPC Cluster of ITP-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the Chroma software suite Edwards and Joo (2005) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020). We also thank Christine Davies for helpful discussion. The numerical calculation were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS and Advanced Computing East China Sub-center. This work is supported in part by NSFC grants No. 12293060, 12293062, 12293061, 12293063, 12293064, 12293065, 12047503, 12175279 and 11935017, the science and education integration young faculty project of University of Chinese Academy of Sciences, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDB34030303, XDB34030301 and YSBR-101, and also a NSFC-DFG joint grant under Grant No. 12061131006 and SCHA 458/22.
Appendix
IV.1 Autocorrelation
@@
along with the masses of pion and $\eta_{c}$. Figs. 13 and 14 show the normalized variance $\sigma_{n}/\sigma_{1}$ of these $4$ quantities as functions of the bin size $n$ within each ensemble.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption

### v3

@@ IV Summary
In this work, we tuned the valence strange quark mass using fictitious $\eta_{s}$ mass Borsanyi et al. (2021) which corresponds to the physical strange quark mass and effectively suppress the impact of the miss-tuning effect strange quark mass of the CLQCD ensembles. We further tune the valence charm quark mass based on the pure QCD $m_{D_{s}}^{\rm QCD}$ for each ensemble, effectively reducing the uncertainty in our prediction stemming from lattice spacing determination, the scale parameter $w_{0}$, and discretization errors in the charm quark mass.
@@
Comparing to the overlap fermion where $Z_{V}=Z_{A}$ only depends on $m_{\rm PS}^{2}a^{2}$ with a tiny coefficient $\sim 0.02$ Wang et al. (2021); He et al. (2022), the $Z_{V}$ of the clover fermion we used here has a much stronger $m_{\rm PS}^{2}a^{2}$ dependence with a coefficient $\sim 0.5$. It makes the heavy quark improved $Z^{c}_{V}$ normalization to be essential here to suppress the discretization error of the charmed meosn matrix elements.
Eventually, we have succeeded in reducing the discretization error of both the charmed hadron mass and the decay constants of open charm hadrons to a few percent level at the coarsest lattice spacing of $a=0.105$ fm. Such a lattice spacing is fairly cost-effective for investigating the dependence of charmed hadrons on light and strange quarks with high precision. As evidence, the mass difference between $D^{+}$ and $D^{0}$ is predicted to be 5.4(5) MeV, incorporating QED corrections from the literature Di Carlo et al. (2019), which aligns well with the experimental value of 4.8(1) MeV. We anticipate that the uncertainty in our prediction can be significantly reduced through direct QED correction calculations on the CLQCD ensemble using infinite volume reconstruction techniques Feng and Jin (2019).
We also verified the unitarily of the CKM matrix elements involving the charm quark at 3% level, updated the constraints on the leptonic decay width of $D^{*}_{(s)}$, and provided the decay constants of the P-wave charmonium decay constants in the continuum with physical quark masses.
However, the hyperfine splittings such as $m_{\rm V}-m_{\rm PS}$ and $\Delta_{1P_{\rm spin-orbit},\bar{c}c}$ still have significant discretiation error, when the charm quark employs the identical discretized fermion action as the light quarks. This effect can hinder the efficient investigation at the coarse lattice spacing, for physics observables which are sensitive to these splittings. Introducing an additional $1+{\cal O}(a^{2})$ parameter in the heavy quark action, as proposed in Liu et al. (2010), could further reduce the discretization error of the hyperfine splitting, as indicated in Brown et al. (2014), and warrants further exploration.
@@
In summary, we would expect the previous Zhang et al. (2022c); Xing et al. (2022); Liu et al. (2023, 2024a, 2024b); Meng et al. (2024a); Han et al. (2024); Yan et al. (2024); Meng et al. (2024b) and ongoing studies using the CLQCD ensembles for the charm physics can reach higher accuracy with the tuned valence strange and charm quark masses and also the heavy quark improved normalization scheme presented in this work. More specifically, hopefully, the systematic uncertainties coming from the unphysical light and strange quark masses, finite volume, and finite lattice spacing can be controlled using the CLQCD ensembles, in the spectrum, form factors, semileptonic decay, and behaviors at the finite temperature and magnetic field of light and charmed meson and baryon.
@@ Acknowledgement
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu et al. (2024), which are generated on HPC Cluster of ITP-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the Chroma software suite Edwards and Joo (2005) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020). We also thank Christine Davies, Simon Kuberski, and Tobias Tsang for helpful discussion and comments. The numerical calculations were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS, and Advanced Computing East China Sub-center. This work is supported in part by NSFC grants No. 12293060, 12293062, 12293061, 12293063, 12293064, 12293065, 12435002, 12047503, 12175279, 11935017, the science and education integration young faculty project of University of Chinese Academy of Sciences, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDB34030303, XDB34030301 and YSBR-101, and also a NSFC-DFG joint grant under Grant No. 12061131006 and SCHA 458/22.
Supplemental Materials
IV.1 Detailed information of CLQCD ensembles and the autocorrelation
Table 3: Gauge coupling $\hat{\beta}=10/(g^{2}u_{0}^{4})$, lattice spacing $a$ with the second error from that of the scale parameter $w_{0}$, tadpole improvement factors $u_{0}$ and $v_{0}$, critical bare quark mass parameters $\tilde{m}^{\mathrm{crti}}$ which drives the pion mass towards zero, dimensionless bare quark mass parameters $\tilde{m}^{\rm b}_{l,s}$, Lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$ and $\eta_{s}$ mass $m_{\eta_{s}}$, valence bare strange and charm quark mass parameters $\tilde{m}^{\rm v}_{s,c}$.
Symbol $\hat{\beta}$ $a$ (fm) $u_{0}$ $v_{0}$ $\tilde{m}^{\mathrm{crti}}$ $\tilde{m}^{b}_{l}$ $\tilde{m}^{b}_{s}$ $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $m_{\eta_{s}}$ (MeV) $\tilde{m}^{\rm v}_{s}$ $\tilde{m}^{\rm v}_{c}$
C24P34 6.200 0.10524(05)(62) 0.855453 0.951479 --0.28854(7) --0.2770 --0.2310 $24^{3}\times 64$ 340.2(1.7) 748.61(75) --0.2396(1) 0.4072(07)
C24P29 0.855453 0.951479 --0.28557(4) --0.2770 --0.2400 $24^{3}\times 72$ 292.3(1.0) 657.83(64) --0.2356(1) 0.4159(07)
C32P29 0.855453 0.951479 --0.28570(3) --0.2770 --0.2400 $32^{3}\times 64$ 293.1(0.8) 658.80(43) --0.2358(1) 0.4150(06)
C32P23 0.855520 0.951545 --0.28432(4) --0.2790 --0.2400 $32^{3}\times 64$ 227.9(1.2) 643.93(45) --0.2337(1) 0.4190(07)
C48P23 0.855520 0.951545 --0.28422(5) --0.2790 --0.2400 $48^{3}\times 96$ 224.1(1.2) 644.08(62) --0.2338(1) 0.4196(08)
C48P14 0.855548 0.951570 --0.28436(2) --0.2825 --0.2310 $48^{3}\times 96$ 136.4(1.7) 706.55(39) --0.2335(1) 0.4205(07)
E28P35 6.308 0.08973(20)(53) 0.859646 0.954385 --0.25871(5) --0.2490 --0.2170 $28^{3}\times 64$ 351.4(1.4) 717.94(93) --0.2201(3) 0.2823(25)
F32P30 6.410 0.07753(03)(45) 0.863437 0.956942 --0.23539(3) --0.2295 --0.2050 $32^{3}\times 96$ 300.4(1.2) 675.98(97) --0.2038(1) 0.1974(05)
F48P30 0.863473 0.956984 --0.23547(2) --0.2295 --0.2050 $48^{3}\times 96$ 302.7(0.9) 674.76(58) --0.2037(1) 0.1965(04)
F32P21 0.863488 0.957017 --0.23477(3) --0.2320 --0.2050 $32^{3}\times 64$ 210.3(2.3) 658.79(94) --0.2023(1) 0.1996(04)
F48P21 0.863499 0.957006 --0.23478(2) --0.2320 --0.2050 $48^{3}\times 96$ 207.5(1.1) 661.94(64) --0.2025(1) 0.1997(04)
G36P29 6.498 0.06887(12)(41) 0.866476 0.958910 --0.21981(2) --0.2150 --0.1926 $36^{3}\times 108$ 297.2(0.9) 693.05(46) --0.1928(1) 0.1433(12)
H48P32 6.720 0.05199(08)(31) 0.873378 0.963137 --0.18884(4) --0.1850 --0.1700 $48^{3}\times 144$ 316.6(1.0) 691.88(65) --0.1701(1) 0.0551(07)
The results in this work, are based on the 2+1 flavor ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action.
The TITLS gauge action, denoted as $S_{g}$, is defined in the following,
$S_{g}=\frac{1}{N_{c}}\mathrm{Re}\sum_{x,\mu<\nu}\mathrm{Tr}\big{[}1-\hat{\beta}\big{(}\mathcal{P}^{U}_{\mu,\nu}(x)+\frac{c_{1}\mathcal{R}^{U}_{\mu,\nu}(x)}{1-8c_{1}^{0}}\big{)}],$ (19)
where $N_{c}=3$, $\mathcal{P}^{U}_{\mu,\nu}(x)=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x)$, $\mathcal{R}^{U}_{\mu,\nu}(x)=U_{\mu}(x)U_{\mu}(x+a\hat{\mu})U_{\nu}(x+2a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\mu}+a\hat{\nu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x)$, $U_{\mu}(x)=P[\mathrm{exp}({\rm i}g_{0}\int_{x+\hat{\mu}a}^{x}\mathrm{d}yA_{\mu}(y))]$, $\hat{\beta}=(1-8c_{1}^{0})\frac{6}{g_{0}^{2}u_{0}^{4}}\equiv 10/(g_{0}^{2}u_{0}^{4})$ with $c_{1}^{0}=-\frac{1}{12}$, $c_{1}=\frac{c_{1}^{0}}{u_{0}^{2}}$, $u_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{U}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$ is the tadpole improvement factor, $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ is the dimensionless 4-D volume of the lattice, and we use $\tilde{O}$ for the dimensionless value of any quantity $O$.
The TITLC fermion action uses 1-step stout smeared link $V$ with smearing parameter $\rho=0.125$,
$S_{q}(m)=\sum_{x,\mu=1,...,4,\eta=\pm}\bar{\psi}(x+\eta\hat{\mu}a)\frac{1-\eta\gamma_{\mu}}{2}V_{\eta\mu}(x)\psi(x)$
$\quad+\sum_{x}\bar{\psi}(x)\big{[}-(4+ma)\delta_{y,x})+c_{\rm sw}\sigma^{\mu\nu}g_{0}F^{V}_{\mu\nu}\big{]}\psi(x),$ (20)
where $c_{\rm sw}=\frac{1}{v^{3}_{0}}$ with $v_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{V}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$, and
$F^{V}_{\mu\nu}$ $=$ $\frac{i}{8a^{2}g_{0}}(\mathcal{P}^{V}_{\mu,\nu}-\mathcal{P}^{V}_{\nu,\mu}+\mathcal{P}^{V}_{\nu,-\mu}-\mathcal{P}^{V}_{-\mu,\nu}$ (21)
$+\mathcal{P}^{V}_{-\mu,-\nu}-\mathcal{P}^{V}_{-\nu,-\mu}+\mathcal{P}^{V}_{-\nu,\mu}-\mathcal{P}^{V}_{\mu,-\nu}).$
Note that the TITLC action here uses the smeared gauge field $V$ for the clover term $\sigma^{\mu\nu}F^{V}_{\mu\nu}$ to suppress the statistical fluctuation, and then different from the action defined in Ref. Horsley et al. (2008); Cundy et al. (2009) which uses $U$ for the clover term.
In addition to the ensembles employed for determining the light quark masses and low energy constants Hu et al. (2024), we have included two additional ensembles, E28P35 and G36P29, with $m_{\pi}\sim 300$ MeV and gauge couplings $\hat{\beta}=6.308$ and 6.498 to manage discretization errors. The tuning of the action parameters at these two values of $\hat{\beta}$ is similar to that in Ref. Hu et al. (2024), and the lattice spacings are determined by the scale parameter $w_{0}$ Aoki et al. (2022); Borsanyi et al. (2012); Bazavov et al. (2014); Blum et al. (2016) which is obtained using the gradient flow Lüscher (2010) of the same gauge action. The spatial volume of the E28P35 and G36P29 ensembles are similar to those of the other ensembles with $m_{\pi}\sim$ 300 MeV. The parameters employed for the simulation are outlined in Table 3.
@@ IV.1 Detailed information of CLQCD ensembles and the autocorrelation
    Refer to caption Refer to caption
    Refer to caption Refer to caption
    Refer to caption Refer to caption
@@
along with the masses of pion and $\eta_{c}$. Figs. 5 and 6 show the normalized variance $\sigma_{n}/\sigma_{1}$ of these $4$ quantities as functions of the bin size $n$ within each ensemble. The mild $n$ dependence of $\sigma_{n}/\sigma_{1}$ suggests that the autocorrelation between configurations is acceptable, even though some of the cases have large statistical uncertainty. This is particularly evident in the case of $Q$, which does not exhibit any indications of topological charge freezing at the finest lattice spacing $a=0.0520(2)(3)$ fm.